# SemCoT: Accelerating Chain-of-Thought Reasoning through Semantically-Aligned Implicit Tokens

### v1

@@ 4.3 Ablation Study
We aim to answer RQ2 by evaluating the effect of different components in SemCoT with its three variants: (1) SemCoT-NST: we remove the customized sentence transformer and modify the semantic alignment loss as the cosine similarity between the mean-pooled embeddings of the implicit reasoning tokens and the ground-truth reasoning; (2) SemCoT-NSA: we remove the entire semantic alignment loss when training the implicit reasoning generator; (3) SemCoT-NLL: we replace the lightweight language model in the implicit reasoning generator with the original LLM and finetune the LLM using LoRA hu2022lora to generate implicit reasoning. We conduct experiments with LLama-2-7B-chat-hf llama2 on various datasets and observe from Fig. 3 that (1) Generally, SemCoT’s performance decrease when any component is removed, showing that each component of the SemCoT positively contributes to the CoT performance; (2) Comparing Fig. 3 (a) and Fig. 3 (b), we find that removing the semantic alignment loss results in worse performance compared to that of optimizing with semantic alignment measured with cosine similarity, highlighting the significance of the semantic alignment loss in learning implicit reasoning; (3) From Fig. 3 (c), we surprisingly find that finetuning the original LLM to generate the implicit reasoning is less effective than utilizing the lightweight LMs, likely attributed to catastrophic forgetting induced by LLM finetuning xu2025softcot .
@@ 4.4 Parameter Sensitivity
We answer RQ3 by studying the fluctuation of SemCoT performance w.r.t. the change of hyper-parameter $\lambda$. Here, $\lambda$ controls the weight of the semantic alignment loss when training the implicit reasoning generator. More specifically, we vary $\lambda$ in $\{0.1,0.3,0.5,0.7,0.9\}$, and we present the corresponding performance differences of SemCoT when applied to Llama-2-7b-chat-hf llama2 on SVAMP patel-etal-2021-nlp . In addition, we also investigate the effect of the number of implicit reasoning tokens $M$ when LLM performs inference. We vary its value among $M=\{1,2,3,4,5\}$ and record the performance of SemCoT using the same setting as when evaluating $\lambda$. The results are shown in Fig. 4. We make the following observations: (1) The LLM answer accuracy generally increases except for when $\lambda$ is 0.5, potentially due to the delicate balance between semantic alignment and prediction accuracy. This is also indicated by the unusual large variance. The inference time is consistent under different $\lambda$. (2) The LLM answer accuracy decreases as the number of implicit tokens increase, indicating that the implicit reasoning can concisely embed ground-truth reasoning information. We observe similar findings in other datasets as well.
@@ 4.5 Case Study
We answer RQ4 through a case study by comparing the implicit reasoning of SemCoT with the baselines. Since directly decoding implicit reasoning to human language is challenging (see Section 3.2), we adopt a proxy evaluation based on this assumption: if an implicit CoT method effectively generates semantically-aligned implicit reasoning from ground-truth reasoning, it should produce consistent implicit reasoning when processing semantically aligned queries. We thus randomly select three queries from the SVAMP patel-etal-2021-nlp dataset and use GPT-4o-mini OpenAI2024GPT4omini to generate 20 semantically equivalent variants for each query. We visualize (with PCA gewers2021principal ) the first implicit reasoning tokens produced by both SemCoT (blue) and the baseline COCONUT (orange) when processing these variants (see case studies for other baseline methods in Appendix D.3). As shown in Fig. 5, the implicit reasoning generated by SemCoT is significantly more concentrated compared to that of COCONUT. The tight clustering suggests that SemCoT successfully distills essential reasoning information regardless of surface-level linguistic differences in queries.
@@ 5 Related Work
Efficient CoT with Implicit Reasoning. Stepwise or progressive encoding methods encode ground-truth reasoning implicitly by first familiarizing the LLM with the reasoning and then removing it, expecting the LLM to learn hidden embeddings that capture previously learned information  deng2024explicit ; hao2024training ; shen2025efficient . These techniques are indirect and susceptible to noise. Other methods compress parts of the reasoning sequence, with CCoT cheng2024compressed focusing on keywords and Token Assorted su2025token selecting parts through a specially trained neural network. SoftCoT xu2025softcot generates the implicit reasoning with a small language model and only trains a linear layer to transform the implicit reasoning of the small LM to the LLM for answer correctness, completely bypassing ground-truth reasoning. CODI shen2025codi employs self-distillation where the teacher model learns ground-truth reasoning while the student learns to generate correct answers with implicit reasoning, the teacher distills the LLM embedding of the last query token to the student in order to encode ground-truth reasoning semantics to the student. This approach lacks verification that the final token fully encodes reasoning semantics. These methods barely consider token-level reasoning generation speed and reasoning semantic alignment.
Semantic Textual Similarity. Semantic Textual Similarity (STS) has evolved from word-level lexical li2006sentence ; mihalcea2006corpus approaches to word-level semantic mikolov2013distributed ; pennington2014glove methods, and finally to sentence-level semantic reimers2019sentence ; cer2018universal ; conneau2017supervised techniques. Early methods relied on lexical overlap and WordNet-based measures, often struggling with synonymy and polysemy li2006sentence . Word-level semantic embeddings like Word2Vec mikolov2013distributed and GloVe pennington2014glove marked a significant advancement, with embeddings typically aggregated to create comparable sentence representations. Recent specialized neural architectures focus on sentence-level representations, including Sentence-BERT (SBERT) reimers2019sentence , which modified BERT with siamese networks to derive semantically meaningful sentence embeddings that can be compared using cosine similarity. Similar approaches include Universal Sentence Encoder cer2018universal and InferSent conneau2017supervised , which share the goal of creating fixed-length sentence representations. Recent works improve embedding space uniformity through contrastive learning strategies gao2021simcse ; zhuo2023whitenedcse ; li2024aoe . However, these STS methods cannot compare semantics between LLMs’ embedding and natural language text.
@@ D.1 Ablation Study
In this section, we show the supplementary experiment results for the ablation study. Specifically, we adopt the variants of the SemCoT as designed in the Section 4.3 in the main paper and examine their performance on all datasets (GSM8K [5], SVAMP [38], MultiArith [4], CommonsenseQA [45], and CoinFlip [22]) on both LLMs (Llama-2-7b-chat-hf [47] and Mistral-7B-Instruct-v0.2 [20]). The results for Llama-2-7b-chat-hf [47] and Mistral-7B-Instruct-v0.2 [20] are shown in the Fig 6 and Fig. 7 respectievly. From the two figures, we observe that our SemCoT performs better than all variants in almost all experiment configurations (i.e., the composition of datasets and LLMs), and the three observations from Section 4.3 also hold.
@@ D.3 Case Study: Semantic Alignment Analysis
We conduct extensive experiments to show that our method helps maintain the semantic alignment between ground truth and implicit reasoning. As declared in Section 4.5, we randomly pick three samples from each dataset and generate semantically aligned queries for each query twenty times. Then we design the experiments to show that our SemCoT achieves semantically-aligned reasoning based on three assumptions: (1) semantically aligned queries will induce semantically aligned reasoning; (2) Different queries and their semantic aligned variants should be well-separated in implicit reasoning space to differentiate their semantics in the implicit reasoning space; (3) Samples from different semantic domains should be more separated than those samples from the same semantic domain to differentiate their semantics domain in the implicit reasoning space. Here, the “semantic domain” means the semantic type of the reasoning task. For example, GSM8K [5], SVAMP [38], MultiArith [4] are for mathematical reasoning, they are in the same semantic domain. However, CommonsenseQA [45] is for commonsense reasoning. Thus, the samples within GSM8K [5], SVAMP [38], and MultiArith [4] are in the same semantic domain, but their samples are in different semantic domains with CommensenseQA. The results are shown in the Fig. 12, Fig. 13, Fig. 14, Fig. 15, Fig. 16, and Fig. 17. For each figure, each subplot of the first row is the implicit reasoning tokens of all samples and their semantic-aligned variants from one dataset. In the second row, each subplot is the $i$th sample of all datasets, along with their semantic-aligned variants. We can observe across the five figures that our SemCoT is the only model to maintain low implicit reasoning variance under semantic-aligned queries and appropriately separate different samples in the implicit reasoning space. Meanwhile, we can also see that our SemCoT recognizes semantic domains because the implicit reasoning for mathematical reasoning is reasonably separate from other domains, such as commonsense reasoning.

### v2

@@ 4.3 Ablation Study
We aim to answer RQ2 by evaluating the effect of different components in SemCoT with its three variants: (1) SemCoT-NST: we remove the customized sentence transformer and modify the semantic alignment loss as the cosine similarity between the mean-pooled embeddings of the implicit reasoning tokens and the ground-truth reasoning; (2) SemCoT-NSA: we remove the entire semantic alignment loss when training the implicit reasoning generator; (3) SemCoT-NLL: we replace the lightweight language model in the implicit reasoning generator with the original LLM and finetune the LLM using LoRA Hu et al. (2022) to generate implicit reasoning. We conduct experiments with LLama-2-7B-chat-hf Touvron et al. (2023) on various datasets and observe from Fig. 3 that (1) Generally, SemCoT’s performance decrease when any component is removed, showing that each component of the SemCoT positively contributes to the CoT performance; (2) Comparing Fig. 3 (a) and Fig. 3 (b), we find that removing the semantic alignment loss results in worse performance compared to that of optimizing with semantic alignment measured with cosine similarity, highlighting the significance of the semantic alignment loss in learning implicit reasoning; (3) From Fig. 3 (c), we surprisingly find that finetuning the original LLM to generate the implicit reasoning is less effective than utilizing the lightweight LMs, likely attributed to catastrophic forgetting induced by LLM finetuning Xu et al. (2025).
@@ 4.4 Parameter Sensitivity
We answer RQ3 by studying the fluctuation of SemCoT performance w.r.t. the change of hyper-parameter $\lambda$. Here, $\lambda$ controls the weight of the semantic alignment loss when training the implicit reasoning generator. More specifically, we vary $\lambda$ in $\{0.1,0.3,0.5,0.7,0.9\}$, and we present the corresponding performance differences of SemCoT when applied to Llama-2-7b-chat-hf Touvron et al. (2023) on SVAMP Patel et al. (2021). In addition, we also investigate the effect of the number of implicit reasoning tokens $M$ when LLM performs inference. We vary its value among $M=\{1,2,3,4,5\}$ and record the performance of SemCoT using the same setting as when evaluating $\lambda$. The results are shown in Fig. 4. We make the following observations: (1) The LLM answer accuracy generally increases except for when $\lambda$ is 0.5, potentially due to the delicate balance between semantic alignment and prediction accuracy. This is also indicated by the unusual large variance. The inference time is consistent under different $\lambda$. (2) The LLM answer accuracy decreases as the number of implicit tokens increase, indicating that the implicit reasoning can concisely embed ground-truth reasoning information. We observe similar findings in other datasets as well.
@@ 4.5 Case Study
We answer RQ4 through a case study by comparing the implicit reasoning of SemCoT with the baselines. Since directly decoding implicit reasoning to human language is challenging (see Section 3.2), we adopt a proxy evaluation based on this assumption: if an implicit CoT method effectively generates semantically-aligned implicit reasoning from ground-truth reasoning, it should produce consistent implicit reasoning when processing semantically aligned queries. We thus randomly select three queries from the SVAMP Patel et al. (2021) dataset and use GPT-4o-mini Menick et al. (2024) to generate 20 semantically equivalent variants for each query. We visualize (with PCA Gewers et al. (2021)) the first implicit reasoning tokens produced by both SemCoT (blue) and the baseline COCONUT (orange) when processing these variants (see case studies for other baseline methods in Appendix D.3). As shown in Fig. 5, the implicit reasoning generated by SemCoT is significantly more concentrated compared to that of COCONUT. The tight clustering suggests that SemCoT successfully distills essential reasoning information regardless of surface-level linguistic differences in queries.
@@ 5 Related Work
Efficient CoT with Implicit Reasoning. Stepwise or progressive encoding methods encode ground-truth reasoning implicitly by first familiarizing the LLM with the reasoning and then removing it, expecting the LLM to learn hidden embeddings that capture previously learned information  Deng et al. (2024); Hao et al. (2024); Shen et al. (2025a). These techniques are indirect and susceptible to noise. Other methods compress parts of the reasoning sequence, with CCoT Cheng and Van Durme (2024) focusing on keywords and Token Assorted Su et al. (2025) selecting parts through a specially trained neural network. SoftCoT Xu et al. (2025) generates the implicit reasoning with a small language model and only trains a linear layer to transform the implicit reasoning of the small LM to the LLM for answer correctness, completely bypassing ground-truth reasoning. CODI Shen et al. (2025b) employs self-distillation where the teacher model learns ground-truth reasoning while the student learns to generate correct answers with implicit reasoning, the teacher distills the LLM embedding of the last query token to the student in order to encode ground-truth reasoning semantics to the student. This approach lacks verification that the final token fully encodes reasoning semantics. These methods barely consider token-level reasoning generation speed and reasoning semantic alignment.
Semantic Textual Similarity. Semantic Textual Similarity (STS) has evolved from word-level lexical Li et al. (2006); Mihalcea et al. (2006) approaches to word-level semantic Mikolov et al. (2013); Pennington et al. (2014) methods, and finally to sentence-level semantic Reimers and Gurevych (2019); Cer et al. (2018); Conneau et al. (2017) techniques. Early methods relied on lexical overlap and WordNet-based measures, often struggling with synonymy and polysemy Li et al. (2006). Word-level semantic embeddings like Word2Vec Mikolov et al. (2013) and GloVe Pennington et al. (2014) marked a significant advancement, with embeddings typically aggregated to create comparable sentence representations. Recent specialized neural architectures focus on sentence-level representations, including Sentence-BERT (SBERT) Reimers and Gurevych (2019), which modified BERT with siamese networks to derive semantically meaningful sentence embeddings that can be compared using cosine similarity. Similar approaches include Universal Sentence Encoder Cer et al. (2018) and InferSent Conneau et al. (2017), which share the goal of creating fixed-length sentence representations. Recent works improve embedding space uniformity through contrastive learning strategies Gao et al. (2021); Zhuo et al. (2023); Li and Li (2024). However, these STS methods cannot compare semantics between LLMs’ embedding and natural language text.
@@ D.1 Ablation Study
In this section, we show the supplementary experiment results for the ablation study. Specifically, we adopt the variants of the SemCoT as designed in the Section 4.3 in the main paper and examine their performance on all datasets (GSM8K [5], SVAMP [36], MultiArith [4, 39], CommonsenseQA [44], and CoinFlip [21]) on both LLMs (Llama-2-7b-chat-hf [46] and Mistral-7B-Instruct-v0.2 [19]). The results for Llama-2-7b-chat-hf [46] and Mistral-7B-Instruct-v0.2 [19] are shown in the Fig 6 and Fig. 7 respectievly. From the two figures, we observe that our SemCoT performs better than all variants in almost all experiment configurations (i.e., the composition of datasets and LLMs), and the three observations from Section 4.3 also hold.
@@ D.3 Case Study: Semantic Alignment Analysis
We conduct extensive experiments to show that our method helps maintain the semantic alignment between ground truth and implicit reasoning. As declared in Section 4.5, we randomly pick three samples from each dataset and generate semantically aligned queries for each query twenty times. Then we design the experiments to show that our SemCoT achieves semantically-aligned reasoning based on three assumptions: (1) semantically aligned queries will induce semantically aligned reasoning; (2) Different queries and their semantic aligned variants should be well-separated in implicit reasoning space to differentiate their semantics in the implicit reasoning space; (3) Samples from different semantic domains should be more separated than those samples from the same semantic domain to differentiate their semantics domain in the implicit reasoning space. Here, the “semantic domain” means the semantic type of the reasoning task. For example, GSM8K [5], SVAMP [36], MultiArith [4, 39] are for mathematical reasoning, they are in the same semantic domain. However, CommonsenseQA [44] is for commonsense reasoning. Thus, the samples within GSM8K [5], SVAMP [36], and MultiArith [4, 39] are in the same semantic domain, but their samples are in different semantic domains with CommensenseQA. The results are shown in the Fig. 12, Fig. 13, Fig. 14, Fig. 15, Fig. 16, and Fig. 17. For each figure, each subplot of the first row is the implicit reasoning tokens of all samples and their semantic-aligned variants from one dataset. In the second row, each subplot is the $i$th sample of all datasets, along with their semantic-aligned variants. We can observe across the five figures that our SemCoT is the only model to maintain low implicit reasoning variance under semantic-aligned queries and appropriately separate different samples in the implicit reasoning space. Meanwhile, we can also see that our SemCoT recognizes semantic domains because the implicit reasoning for mathematical reasoning is reasonably separate from other domains, such as commonsense reasoning.